OpenStudy (anonymous):

What is the coefficient of x^9 in (2+x)12?

OpenStudy (anonymous):

And you ment to raise the ^12

OpenStudy (anonymous):

I*

OpenStudy (anonymous):

General term is \[T _{r+1}=_{_{r}^{C}}^{12}\left( 2 \right)^{12-r}\left( x \right)^r\] put r=9 the coefficient of x^9 is \[C_{9}^{12}\left( 2 \right)^{12-9}=?\]

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

That was my first answer, I don;t see what I'm doing wrong.

OpenStudy (anonymous):

don't

OpenStudy (anonymous):

\[C_{9}^{12}2^{12-9}=C _{12-9}^{12}2^{3}=C _{3}^{12}(8)=\frac{ 12*11*10 }{ 3*2*1 }*8=?\]

OpenStudy (anonymous):

ohhhhh, so you actually multiple the combination part! Sweet! Thank you!